The shift was dedicated to the planned ISC activity, LN1 Recovery, carried out by Casanueva and De Rossi. Unfortunately, the foreseen work could not be performed due to the worsening weather condition, with strong winds and rain (see attached plots). The crew also noted an incorrect behavior of the PR ChRoCC temperature, investigation in progress.
OSD Calibrations
7:04 UTC - 5 minutes of Single bounce NI;
7:09 UTC - 5 minutes of Single bounce WI;
7:33 UTC - 5 minutes of CITF lock.
SUSP
From 10:42 UTC to 10:56 UTC, Ruggi reduced the value of Sa_WI_f0_COIL_H2.
TCS
To study the behavior of the PR CHROCC, under request of the crew I changed its V DAC Value. Here's the timetable of the switch performed:
9:42 UTC - from 0.38 V to 0.4 V;
10:12 UTC - from 0.4 V to 1.5 V;
10:38 UTC - from 1.5 V to 0.38 V.